If 456 dm3

of krypton at 101 kPa and 21°C is compressed into a 27.0 dm3

tank at the

same temperature, what is the pressure of krypton in the tank?


1
Expert's answer
2021-05-31T02:06:56-0400

As at same temperature , P1 × V1 = P2 × V2


Where P1 and P2 are initial and final pressures respectively and V1 and V2 are initial and final volumes respectively;


Now , P1 = 101 kPa , V1 = 456 dm3 , V2 = 27.0 dm3


So , P2 = ( P1 × V1 ) / V2


= ( 101 kPa × 456 dm3 ) / 27.0 dm3


= 1705.7778 kPa